5x + 2y = 8
x + y = 4
If you want to solve the system of equations by addition, which of the following could you do?

Answer by stanbon(75887)

5x + 2y = 8
x + y = 4
If you want to solve the system of equations by addition
----
Multiply thru the bottom equation by -5::
5x + 2y = 8
-5x-5y = -20
---
Add and solve for "y":
-3y = -12
----
Divide by -3 to get:
y = 4
---
Solve for "x":
x + y = 4
x + 4 = 4
x = 0
================
Cheers,
Stan H.
==================